Estimate the peak wavelength for radiation from the following sources, assuming blackbody emission.

(a) ice at -4
Answer #1

apply Weins dispalcement law LT = 0.0029 mK

so

--------------------------------------------------------------------------------

a. L = 0.0029/(-4+273) = 10.78 um

falls in 3. infrared, microwave, near infrared, and far infrared respectively

--------------------------------------------------------------------------------

b. L = 0.0029/2900 = 1000 nm

2. far infrared, near infrared, microwave, and infrared respectively    

---------------------------------------------------------------------------------

c. L = 0.0029/4   = 0.725 mm

5. infrared, microwave, far infrared, and near infrared respectively

------------------------------------------------------------------------------------

d. L = 0.0029/2.725 = 0.106 cm

5. infrared, microwave, far infrared, and near infrared respectively

--------------------------------------------------------------